OpenStudy (anonymous):

factor the polynomial 20w^7+28w^3

OpenStudy (campbell_st):

well 4w^3 seems like a common factor

OpenStudy (anonymous):

where did the 3 come from

OpenStudy (campbell_st):

its \[4w^3\]

OpenStudy (anonymous):

i mean the exponet 3

OpenStudy (campbell_st):

well \[w^7 = w^4 \times w^3\]

OpenStudy (campbell_st):

so the question can be written as \[20w^4 \times w^3 + 28w^3\]

OpenStudy (campbell_st):

just identify the common factors